OpenStudy (anonymous):

If the tangent line to y = f(x) at (4, 2) passes through the point (0, 1), find f(4) and f '(4). f(4)=2 , f'(4)=?

myininaya (myininaya):

so you can find the tangent line you are given two points and the slope use point-slope form of a line

OpenStudy (anonymous):

Y1-Y0=m(x1-xo)?

myininaya (myininaya):

actually you only need the slope

OpenStudy (anonymous):

yeah

myininaya (myininaya):

since that is what it asked for

myininaya (myininaya):

you know the slope formula right?

myininaya (myininaya):

\[m=\frac{y_1-y_2}{x_1-x_2}\]

OpenStudy (anonymous):

m=y1-y0/x1-xo

OpenStudy (anonymous):

is the other way around of how you put it

myininaya (myininaya):

1 and 0's 1 and 2's whatever

myininaya (myininaya):

do you know how to use the formula?

OpenStudy (anonymous):

is 1/4

myininaya (myininaya):

let's see top is 2-1 and bottom is 4-0 looks great

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

do I use y=mx+b?

OpenStudy (anonymous):

b=2 if I do that with (4,2)

myininaya (myininaya):

you don't need to it just asked to find f'(4) and f(4) and f'(4) is the slope of the tangent line at (4,2) to the curve y=f(x)

myininaya (myininaya):

unless you want to find the tangent line ?

OpenStudy (anonymous):

I do

myininaya (myininaya):

it is just: \[y-f(a)=f'(a)(x-a) \text{ is the tangent line \to the curve } y=f(x) \text{ at } (a,f(a))\]

myininaya (myininaya):

you are given a is 4 here

myininaya (myininaya):

\[y-f(4)=f'(4)(x-4)\]

myininaya (myininaya):

now you found f(4) and f'(4)

OpenStudy (anonymous):

right

OpenStudy (anonymous):

thanks

myininaya (myininaya):

np

OpenStudy (anonymous):

problem the answer is not 2.

OpenStudy (anonymous):

I already tried that

myininaya (myininaya):

f(4) should be 2 it is given by (4,2) f'(4) should be 1/4 that is the slope of the tangent line at (4,2) to the curve y=f(x)

OpenStudy (anonymous):

oh yeah then I was right about the 1/4. in the question I put the f(4)

OpenStudy (anonymous):

perfect thank you

myininaya (myininaya):

yeah you already found f(4) earlier we were finding f'(4) and we used the slope formula because that is what the derivative means

OpenStudy (anonymous):

Find f '(a). f(x) = (1 − 2x)^1/2

myininaya (myininaya):

so do you know chain rule?

OpenStudy (anonymous):

yeah, saw this last year tho

OpenStudy (anonymous):

I thought it was 1/2(1-2x)^-1/2 * (2)

myininaya (myininaya):

\[h(x)=f(g(x)) \\ h'(x)=g'(x) \cdot f'(g(x)) \\ \\ \text{ if } f(x)=x^n \\ \text{ then we have } \\ h(x)=(g(x))^n \\ h'(x)=g'(x) \cdot n(g(x))^{n-1}\]

myininaya (myininaya):

well one complaint

myininaya (myininaya):

what is the derivative of 1-2x ?

myininaya (myininaya):

you put 2 but it should actually be...

OpenStudy (anonymous):

ooooh

OpenStudy (anonymous):

thanks

myininaya (myininaya):

but you can also simplify a bit

OpenStudy (anonymous):

hahaha hate when that happens

OpenStudy (anonymous):

yes i did before

OpenStudy (anonymous):

in my homework

myininaya (myininaya):

ok then you got it now :)

OpenStudy (anonymous):

do you have time for more?

OpenStudy (anonymous):

I had problems with this one too

myininaya (myininaya):

maybe one more then I have to go check on something

OpenStudy (anonymous):

(a) Find the slope m of the tangent to the curve y = 7/ x at the point where x = a > 0.

myininaya (myininaya):

is that just y=7/x?

OpenStudy (anonymous):

(b) Find equations of the tangent lines at the points (1, 7) and (4, 7/2) .

myininaya (myininaya):

\[y=\frac{7}{x}=7 \cdot \frac{1}{x} =7 x^{-1} \text{ by law of exponents } \\ \text{ now differentiate } y=7 x^{-1} \text{ using power rule }\]

OpenStudy (anonymous):

no x^1/2 sorry

myininaya (myininaya):

oh

OpenStudy (anonymous):

it didn't appear when i copy pasted it

OpenStudy (anonymous):

I had -7/2x^3/2

OpenStudy (anonymous):

but is not that

myininaya (myininaya):

\[\text{ find the slope of the tangent line \to the curve } y=7x^\frac{1}{2} \text{ at } x=a>0\]?

myininaya (myininaya):

or is that x^(1/2) in the denominator?

OpenStudy (anonymous):

ok part A i got it

OpenStudy (anonymous):

I had to change x=a

OpenStudy (anonymous):

now for B I don't now what I did wrong

myininaya (myininaya):

ok so it was: \[y=\frac{7}{x^\frac{1}{2}}=7x^\frac{-1}{2}\]

OpenStudy (anonymous):

I got that

OpenStudy (anonymous):

but I have part B wrong

myininaya (myininaya):

ok the tangent line at (1,7) is \[y-f(a)=f'(a)(x-a) \\ \text{ here you have } a=1 \\ y-f(1)=f'(1)(x-1) \\ \text{ you already found } f'(x) \text{ \above } \\ \text{ you got that } f'(x)=\frac{-7}{2x^{\frac{3}{2}}} \text{ I think but I don't really know anymore } \\ \text{ \because I'm still trying \to figure out what } f(x) \text{ was :p } \\ \text{ so this becomes } \\ y-7=\frac{-7}{2(1)^\frac{3}{2}}(x-1)\]

myininaya (myininaya):

and you still have to do this same process for the other point

OpenStudy (anonymous):

perfect

OpenStudy (anonymous):

is not that apparently